### v2

# font=itshape, raggedright, noindent, leftmargin=0.08in, rightmargin=0.08in, begintext=“ , endtext= ”

Abstract
There is a growing need for investigating how machine learning models operate. With this work, we aim to understand trained machine learning models by questioning their data preferences. We propose a mathematical framework that allows us to probe trained models and identify their preferred samples in various scenarios including prediction-risky, parameter-sensitive, or model-contrastive samples. To showcase our framework, we pose these queries to a range of models trained on a range of classification and regression tasks, and receive answers in the form of generated data.
1 Introduction
Machine learning models are widely used in today’s data-driven world, powering critical decision-making processes in sectors ranging from healthcare to human resources. Their widespread adoption in high-stakes scenarios raises important questions on aligning trained models with human values. Understanding how these models operate has become a critical concern. Our quest along this line starts with the following inquiry: What kind of data can we generate to probe our trained models?
To respond to this inquiry, we study the implicit data distribution favored by trained models. In other words, our approach to understanding a model is based on creating samples in the data domain that the trained model considers favorable for a specific task. Unlike conventional ML pipelines that focus on static datasets and predictive accuracy, our approach enables dynamic interrogation of model behavior via investigating the answers of the model to specific questions. We show that these questions can be customized to each situation and they can be expressed mathematically through a loss function that evaluates the data based on a combination of data characteristics and model parameters. We consider the problem of understanding a model to be a more nuanced endeavor that requires exploration across multiple dimensions of questioning. This involves providing explanations, such as counterfactual [28] or prototypical [3] scenarios, shedding light not only on why a particular prediction was made, but going beyond it as well. For instance, insights into model behavior can be gained by generating parameter-sensitive data samples. When two models showing similar performances give different predictions, also known as predictive multiplicity [22], our approach can be used to generate data to systematically compare their behaviors in diverse scenarios. These custom questions, and others, provide a qualitative understanding of the model. In addition, users have the flexibility to customize queries by designing specific probing functions within the data space.
Related Literature. Our work complements extensive research in synthetic data generation that has been pivotal in addressing fairness, bias detection [15] and reduction [30, 27] as well as dataset augmentation [29, 6]. Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) have been widely used to approximate original data distributions [11, 30, 14, 4], focusing on privacy, diversity, and fidelity as primary goals.
Recent studies leveraged generative models for counterfactual generation and exploring underrepresented data regions. For example, Joshi et al. [12] proposed a framework for generating task-specific synthetic data, enhancing model explainability. Similarly, Redelmeier et al. [25] introduced an approach using autoregressive generative models to create counterfactuals, facilitating bias exploration and decision boundary analysis. Recent work on global counterfactual explanations has further expanded the scope of interpretability by targeting group-level understanding. Rawal and Lakkaraju [24] introduced a framework for generating global, rule-based recourse summaries for population subgroups, optimizing objectives like accuracy, coverage, and cost. These summaries are ‘if-then’ rules linking subgroup characteristics to actionable feature changes that influence a model’s prediction. Plumb et al. [23] proposed interpreting clusters in low-dimensional representations by finding sparse transformations that align one group with another.
Energy-based models (EBMs) have also emerged as a promising framework, combining generative and discriminative modeling tasks. By treating classifier logits as an energy function, EBMs can model joint distributions over data and labels [17, 5]. Applications of EBMs include adversarial robustness, out-of-distribution detection, and data augmentation [31, 19, 1, 21]. For instance, Duvenaud et al. [5] demonstrated improved out-of-distribution detection using a joint energy-based model, while Ma et al. [20] extended EBMs to tabular data for synthetic data generation.
The proposed framework draws inspiration from these works while introducing a distinct perspective. Our probing function can be seen as an energy function and leads to Gibbs distribution. However, rather than learning the energy function to capture the data distribution (conditioned on label), we create a probing function using trained models. This design allows the distribution to generate samples that address the specific posed question. Related works, such as [5] and [20], adopt a similar approach by utilizing a trained classifier to obtain an energy function and using Langevin dynamics for sampling from the Gibbs distribution. However, their main objective is to mimic the true data distribution. In fact, the former paper combines training of the energy function and classifier. In contrast, we propose a flexible framework that allows for directing diverse queries to trained models via probing functions that reflect various objectives, such as identifying prediction-risky, parameter-sensitive, or model-contrastive data samples.
Contributions. We contribute to the literature by introducing a novel inductive approach. This approach creates data samples using a probing function that engages a trained model. Through this framework, the generated samples provide answers to various questions posed to trained models. With a computational study, we support our approach by applying it to a range of classification and regression tasks, showcasing its effectiveness in generating data tailored to specific queries. Our numerical results demonstrate the flexibility of our approach in uncovering biases, facilitating model interpretability, and consequently, promoting alignment of model predictions and human preferences.
2 The Mathematical Framework
First, our notation: The labeled data lie in $\mathcal{X}\times\mathcal{Y}$, and the model defines a predictor function $f({\boldsymbol{\mathbf{\theta}}},\cdot):\mathcal{X}\to\mathcal{Y}^{\prime}$ for any given set of model parameters ${\boldsymbol{\mathbf{\theta}}}\in\Theta$. For a given sample ${\boldsymbol{\mathbf{x}}}\in\mathcal{X}$, the predicted label $y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}})\in\mathcal{Y}$ is obtained from the predictor function. The cost function $\ell_{F}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ measures how far the predicted labels are from the true labels.
Refer to caption Refer to caption
Figure 1: (left) Overview of model probing by data generation. Samples drawn from $p^{*}({\boldsymbol{\mathbf{x}}})$ constitute an answer to the question posed by $G$. The vertical arrows (a) and (c) start with functions and lead to distributions on the same space by solving the Bayesian Learning Problems (2) and (3). The diagonal arrow (b) obtains a loss function on the data space by integrating out the ${\boldsymbol{\mathbf{\theta}}}$ dependence from a designed function via $q^{*}({\boldsymbol{\mathbf{\theta}}})$. (right) Specializing to a linear regression model with mean square error and no regularizer admits an analytic solution. The distribution of data points ${\boldsymbol{\mathbf{x}}}$ which are preferred by the LR model in giving a chosen prediction $y^{\prime}$ is calculated to be a Gaussian distribution centered at a point $\hat{{\boldsymbol{\mathbf{f}}}}$, which is shifted from the mean of given data by a certain amount depending on the desired output value $y^{\prime}$. Explicit forms of $\hat{{\boldsymbol{\mathbf{f}}}},\Sigma$ and the derivation can be seen in Appendix A.
The standard construction of the parameter loss function is
$F({\boldsymbol{\mathbf{\theta}}})=\int_{\mathcal{X}\times\mathcal{Y}}(\ell_{F}(y_{\boldsymbol{\theta}}(\mathbf{x}),y)+R_{F}(\boldsymbol{\theta}))\mathrm{d}\nu(\mathbf{x},y)=\frac{1}{N}\sum_{i=1}^{N}\ell_{F}(y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}_{i}),y_{i})+R_{F}({\boldsymbol{\mathbf{\theta}}}),$ (1)
which can be seen as an integral of $\ell_{F}+R_{F}$ against the empirical distribution given by the training dataset $\{({\boldsymbol{\mathbf{x}}}_{i},y_{i})\}_{i=1}^{N}\subseteq\mathcal{X}\times\mathcal{Y}$. Here, $R_{F}({\boldsymbol{\mathbf{\theta}}})$ is a regularizer term that depends only on ${\boldsymbol{\mathbf{\theta}}}$.
We propose a framework for probing a model with answers in the form of generated data and it is structured to be symmetric to the training process itself: instead of generating parameters given a data distribution, we generate data given a trained model parameter distribution. Similar to how we form the loss $F$ on the parameters, our data loss $G$ can be obtained by integrating out the ${\boldsymbol{\mathbf{\theta}}}$ dependence of a function that depends on both the data and the model parameters. We design this latter function to be low when the desiderata required by a probing scenario are satisfied. Depending on our choice, we end up probing the model with different questions and gain valuable insights into the trained model’s behavior by observing the generated data, both qualitatively and via its population statistics. Figure 1 presents the overview of our framework.
The blue arrows (a) and (c) in Figure 1 represent solving the Bayesian Learning Problem (BLP), which –instead of finding a single set of model parameters ${\boldsymbol{\mathbf{\theta}}}^{*}$– seeks a distribution that balances the tasks of minimizing the expected loss and maximizing the entropy. In case of (a), we solve
$\operatorname*{arg\,min}_{q\in\mathcal{Q}}\mathbb{E}_{q}[F]-\tau\mathcal{H}(q),$ (2)
where $\mathcal{Q}$ is a choice of candidate distributions on $\Theta$, and $\mathcal{H}(q)=-\int_{\Theta}q\log q\,\mathrm{d}\mu$ is the entropy with respect to a base measure $\mu$. The problem can be interpreted as an implementation of the exploration-exploitation trade-off in the parameter space. The constant $\tau>0$, called the temperature, controls the balance between these two objectives. Absent any restriction, i.e., if $\mathcal{Q}$ is the set of all probability distributions [^1] All probability distributions which are absolutely continuous with respect to a given base measure, which in this case taken to be the Lebesgue measure $\mathrm{d}\,{\boldsymbol{\mathbf{\theta}}}$., then the Gibbs-Boltzmann distribution $q^{*}({\boldsymbol{\mathbf{\theta}}})\propto e^{-\frac{1}{\tau}F({\boldsymbol{\mathbf{\theta}}})}$ is the unique solution to (2).
Standard stochastic gradient descent can be seen as a specialization of this setup. Given a very restrictive family $\mathcal{Q}$, such as the manifold of Dirac delta distributions [^2] To eschew technicalities of infinities, instead of exactly using Dirac delta we may instead consider distributions which are supported everywhere, and highly concentrated around a point but with fixed variance. supported on a single ${\boldsymbol{\mathbf{\theta}}}$, the entropy term becomes irrelevant, and we get the classical optimization problem of minimizing $F$.
Symmetrically on the data space $\mathcal{X}$, the blue arrow labeled (c) in Figure 1, represents solving
$\operatorname*{arg\,min}_{p\in\mathcal{P}}\mathbb{E}_{p}[G]-\tau\mathcal{H}(p).$ (3)
The expectation term encourages the solution $p^{*}({\boldsymbol{\mathbf{x}}})$ to concentrate its mass on data samples where $G$ is minimized, but the entropy term encourages $p^{*}({\boldsymbol{\mathbf{x}}})$ to explore a wide variety of data samples. Just like $F$ in (1), the function $G$ is constructed as an integral of a function of two ${\boldsymbol{\mathbf{\theta}}}$ and data against the distribution $q^{*}({\boldsymbol{\mathbf{\theta}}})$. Training in the parameter space encourages finding those parameters where $F$ is lower, so the reached parameters are more compatible with the given training data. Similarly, training with the loss function $G$ in the data space with (3) encourages finding data points which are compatible with the trained parameters sampled from $q^{*}$ (which may be a single set of parameters $\theta^{*}$ in case it is a Dirac-delta distribution). In this way, the samples from the distribution $p^{*}({\boldsymbol{\mathbf{x}}})$ are considered to be an answer to the question put forth by $G$: Which data points makes the $G$-loss low while maintaining diversity among the produced samples? Then, it remains to design $G$ so that this coincides with illuminating questions about the trained model.
We may explicitly select $\mathcal{P}$ as a family of distributions depending on the nature of the data distribution. Alternatively, we can retain the full space of probability distributions but replace $G$ with $G\circ\varphi$ for some function $\varphi:\mathcal{Z}\to\mathcal{X}\times\mathcal{Y}$. In the latter case, $\mathcal{Z}$ can be the latent space and $\varphi$ can be the decoder function of a trained variational autoencoder. This procedure produces a distribution on $\mathcal{Z}$, which we can sample from and map to $\mathcal{X}$ via $\varphi$, thus sampling data from the pushforward distribution.
There are various methods of solving (3). In the examples given below, we used Metropolis Adjusted Langevin Algorithm (MALA) to sample directly from the Gibbs-Boltzmann distribution $p^{*}({\boldsymbol{\mathbf{x}}})\propto e^{{-\frac{1}{\tau}G({\boldsymbol{\mathbf{x}}})}}$. Details of this method are given in Appendix C. Instead of MALA, one can also apply Picard Iterations to solve the overdamped Langevin stochastic differential equation, or alternatively apply Variational Inference (VI) to reach a distribution $p^{*}$; see [8, 9].
3 Probing Trained Models
We start with a general structure of the loss function for questioning trained models:
$G({\boldsymbol{\mathbf{x}}})=\int_{\Theta}\ell_{G}(y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}),\widehat{y}({\boldsymbol{\mathbf{x}}}))q^{*}({\boldsymbol{\mathbf{\theta}}})\mathrm{d}\,{\boldsymbol{\mathbf{\theta}}}+R_{G}({\boldsymbol{\mathbf{x}}}),$ (4)
where $\widehat{y}$ stands for a predictor and $R_{G}$ is a regularizer function that can be chosen to put additional soft constraints on the samples in addition to the hard constraints coming from the restriction $p\in\mathcal{P}$. This general probing function enables us to express a wide range of model-inspection tasks. In the remainder of this section, we walk through several representative cases, each capturing a specific type of question one might ask about a trained model’s behavior. These are not exhaustive, but illustrative scenarios that demonstrate the flexibility of the proposed framework. Figure 2 showcases what sort of data points would be produced in a synthetic dataset.
Figure 2: Given a dataset of two concentric circles labeled red and blue, two Support Vector Machine (SVM) models are trained on the binary classification task with kernels chosen as Radial Basis Function (RBF) and cubic polynomial, respectively. The generated data points are green. In (a), we contrast the two SVM models looking for samples for which their predictions differ and discover that this is the case in a region near the origin lacking any training points. In (b) and (c), we inquire about data points that would be considered risky by the two models using RBF and cubic kernels, respectively. In (d) we start from a data sample shown in orange, and design $G$ so that it generates data points which are classified with the opposite label according to the RBF-SVM without straying too far from the original point.
Model-contrasting samples. Given two models, finding data points where their predictions disagree is illuminating either to interpret model-specific biases or to audit consistency between two different models. This is particularly useful when comparing models with different inductive biases (MLP vs. CNN, linear vs. nonlinear, and so on). Given $\widehat{y}$ as the predictor functions of the model that is being compared against (which can be non-parametric like in XGBoost), we solve (3) with the function
$G({\boldsymbol{\mathbf{x}}})=\ell_{G}(y_{{\boldsymbol{\mathbf{\theta^{*}}}}}({\boldsymbol{\mathbf{x}}}),1-\widehat{y}({\boldsymbol{\mathbf{x}}}))+R_{G}({\boldsymbol{\mathbf{x}}}).$ (5)
In Figure 2(a) we contrast two SVM models with different kernels and discover a region near the origin, which does not contain any samples from the dataset but would give conflicting predictions if a new sample were to come from it.
Prediction-risky samples. In order to identify the indecisive regions in a model’s decision surface, we put $G({\boldsymbol{\mathbf{x}}})=\|f({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{\theta}}}^{*})-{\boldsymbol{\mathbf{\alpha}}}\|_{r}^{r}$ for $r\geq 1$, where $f$ may be the decision function of a model such as in SVM and logistic regression, or it may be the softmax probabilities of a neural network. By solving (3), we thus seek data points with probabilities close to ${\boldsymbol{\mathbf{\alpha}}}$. For binary classification, where $\alpha$ denotes the cut-off point, this corresponds to generating a “risky data point” near the decision boundary. This probing strategy can yield insights into whether the model correctly identifies important aspects of the data for decision-making and whether that aligns with the user’s conceptions. As an illustrative example in Figure 2(b), we search for the decision boundary of a RBF-SVM trained on a dataset of two concentric circles, and we discover the ring in the center as an unexpected decision boundary.
Parameter-sensitive samples. Given a set of parameters ${\boldsymbol{\mathbf{\theta}}}^{*}$ and a distribution $q^{*}({\boldsymbol{\mathbf{\theta}}})$ of parameter values, we ask the model for data samples whose classifications would flip if the model parameters were to be (perhaps slightly) perturbed. This can be achieved by solving (3) using the function
$G({\boldsymbol{\mathbf{x}}})=\int_{\Theta}\ell_{G}(y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}),1-y_{{\boldsymbol{\mathbf{\theta}}}^{*}}({\boldsymbol{\mathbf{x}}}))q^{*}({\boldsymbol{\mathbf{\theta}}})\mathrm{d}\,{\boldsymbol{\mathbf{\theta}}}+R_{G}({\boldsymbol{\mathbf{x}}}).$ (6)
This integral would be approximated by samples from $q^{*}$. When $q^{*}\in\mathcal{Q}$ is chosen from a restricted family of distributions, like Gaussians with fixed variance, sampling from $q^{*}$ means perturbing ${\boldsymbol{\mathbf{\theta}}}^{*}$. This probing function is particularly useful when we want to examine the consistency of a model’s predictions under small shifts in its parameters. By identifying inputs whose predictions vary significantly with minor parameter changes, we can highlight sensitive regions in the input space—areas that might indicate over-dependence on specific parameter configurations. This can have overlaps but is distinct from prediction-risky samples, as we showcase in our computational study section. Parameter-sensitive samples has the flexibility to be generated far from the decision boundary, especially in non-linear models.
Fixed-label samples. Finally, we probe the model for what it thinks are good data samples that fit the bill for the prediction $y^{\prime}$, either for a single parameter ${\boldsymbol{\mathbf{\theta}}}^{*}$ or a distribution $q^{*}({\boldsymbol{\mathbf{\theta}}})$
$G({\boldsymbol{\mathbf{x}}})=\ell_{G}(y_{{\boldsymbol{\mathbf{\theta}}}^{*}}({\boldsymbol{\mathbf{x}}}),y^{\prime})+R_{G}({\boldsymbol{\mathbf{x}}})\quad\text{ and }\quad G({\boldsymbol{\mathbf{x}}})=\int_{\Theta}\ell_{G}(y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}),y^{\prime})q^{*}({\boldsymbol{\mathbf{\theta}}})\mathrm{d}\,{\boldsymbol{\mathbf{\theta}}}+R_{G}({\boldsymbol{\mathbf{x}}}),$ (7)
respectively. Here, $R_{G}({\boldsymbol{\mathbf{x}}})$ is a localizer at an anchor point. We can take a data point $({\boldsymbol{\mathbf{x}}}_{0}.y_{0})$ to be this anchor. In case $y^{\prime}\neq y_{0}$, we are exploring changes in ${\boldsymbol{\mathbf{x}}}_{0}$ that would need to happen for the prediction to change; in other words, a counterfactual; see Figure 2(d). In case $y^{\prime}=y_{0}$ with a weak localizer, we can obtain a sample that would lead to a similar prediction, i.e., a factual.
Figure 1 demonstrates the steps when $y_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}})={\boldsymbol{\mathbf{x}}}^{\intercal}{\boldsymbol{\mathbf{\theta}}}$ corresponds to linear regression, and both $\ell_{F}$ and $\ell_{G}$ are the mean squared errors. For this special case, we obtain analytical solutions for all steps of our framework. The details of this observation are given in Appendix A.
Localized samples. In all of the above cases, we can add a regularizer term $R_{G}({\boldsymbol{\mathbf{x}}})=\|{\boldsymbol{\mathbf{x}}}-{\boldsymbol{\mathbf{x}}}_{a}\|_{r}^{r}$ for $r\geq 1$ to generate synthetic data that is similar locally to an anchor ${\boldsymbol{\mathbf{x}}}_{a}$. In fact, different weightings can also be applied to different columns to enforce this more or less stringently for different features.
[…]
4 Computational Study
In this section, we conduct a series of experiments to evaluate the cases presented in Section 3. Our experiments aim to evaluate the proposed framework by demonstrating its ability to generate data samples across various scenarios. We use well-established datasets that have been recently adopted in related literature (e.g., [10], [18], [26]), and their specifics are outlined in Appendix B. The implementation details and code for reproducing these experiments are available on our GitHub repository. [^3] https://anonymous.4open.science/r/EvD-6FB1/
Model-contrasting samples. This experiment investigates the differences between two predictive models by probing the features that drive contrasting predictions for the same data. Through our framework, we pose the following question:
Which specific features or input changes lead to disagreement between the two models’ predictions?
To explore this, we apply the framework to datasets of different modalities. For tabular data, we use the FICO dataset [7]; for image dataset, we use MNIST [16].
We begin by investigating model divergence in scenarios where the comparison model is non-differentiable. To this end, we train XGBoost -a non-parametric model- alongside logistic regression on the FICO dataset, which consists of credit applications with features related to financial history and risk performance. This setup highlights the flexibility of our framework, as it enables probing differences between models with fundamentally distinct modeling approaches. Although the two models agree on 94.5% of the predictions in the test set, our framework generates a set of samples where their predictions exhibit full disagreement, i.e., XGBoost predicts one class, while logistic regression predicts the opposite. Figure 3 presents the feature distributions for these discrepant samples, focusing on four representative features. We observe that disagreement tends to occur when the number of credit accounts opened in the past year falls in the 12–15 range, which is outside of this feature’s distribution in the test data. This indicates that our method can surface disagreement patterns that would remain undetected through standard evaluation alone. We also include in Appendix D a comparison between a linear model and a Support Vector Regression (SVR) model using a different tabular dataset.
Figure 3: The distributions of four representative features in the generated samples. Here, XGBoost predicts “Bad” for RiskPerformance, while logistic regression predicts “Good”.
Our framework can also be used to compare and contrast two models trained on image data. To demonstrate, we consider a Convolutional Neural Network (CNN) and an MLP, both trained on MNIST. The architectures of these networks are provided in Appendix B.2. To better capture the data manifold, we also train a VAE with a latent dimension of 10. The trained encoder module of the VAE is denoted by $z\mapsto\varphi(z)$. Further details on the VAE training process are provided in Appendix E. In Figure 4, we present an example computation illustrating how this setup works. Starting with a latent vector encoding an image with label ‘3’, we sample from a distribution that prefers the label ‘8’ jointly for both a trained CNN (LeNet5) and an MLP.
We use this setup to systematically compare the CNN and MLP models. In Figure 4, we showcase some samples generated by forcing functions $G$ that pull the data toward incompatible directions, for example, resulting in amorphous data points that exhibit characteristics of both ‘1’ and ‘0’. The third column highlights cases where the label ‘8’ is preferred (top: MLP, bottom: CNN) while remaining close to an actual MNIST image labeled ‘3’, which is enforced through two-norm regularization.
Figure 4: (left) Using Langevin dynamics in the latent space, we obtain a sequence of latent vectors that, when passed through the decoder $\varphi$, correspond to a walk on the data manifold. In this image, the function $G$ is the sum of cross entropy predictions of trained MLP and LeNet5 networks for the label ‘8’ and for the data $\varphi(z)$. (right) Images in the first and second columns are generated to prefer a given label on an MLP model and another one on a CNN model. upper-left: CNN-‘0’ MLP-‘1’, upper-middle: CNN-‘1’ MLP-‘7’, lower-left: CNN-‘0’ MLP-‘8’, lower-middle: CNN-‘2’ MLP-‘5’. On the third column, the upper image prefers the label ‘8’ for the MLP model whilst being close to a data sample with label ‘3’, and the same for the lower image for the CNN model.
Prediction-risky samples. This experiment explores data samples near the model’s decision boundary, where predictions are inherently uncertain. To guide this analysis, we pose the question:
Which data samples are predicted to be risky due to being close to a specific anchor value?
In this experiment, we train a neural network (MLP) to classify customers in the FICO dataset as either “Good” or “Bad” credit risks. Prediction-risky samples are those with model outputs near the anchor value of 0.5. Using our framework, we generate 500 such samples to examine the characteristics of borderline classification cases. The average predicted probability of the “Bad” credit class among these samples is 0.525, with a standard deviation of 0.017.
The density plots in Figure 5 compare the distributions of two representative features in the original data and the generated samples. The feature NumTrades60Ever2DerogPubRec represents the number of past credit trades with payments delayed by at least 60 days, serving as a key indicator of delinquency. As shown in the figure, the distribution of risky samples follows the original data closely in the lower range but exhibits a stronger peak around zero. This suggests that the model considers individuals with few or no past delinquencies as borderline cases, likely due to the absence of strong negative or positive indicators, making classification more uncertain. The feature MSinceOldestTradeOpen indicates the number of months since a customer’s first credit line was opened, capturing the length of their credit history. As shown in Figure 5, the distribution of risky samples is sharply concentrated around 400 months ($\sim$33 years), whereas the original data spans a broader range. This suggests that the model associates long credit histories with greater uncertainty. The pronounced peak reflects how the model treats long-established credit profiles as ambiguous when making predictions. The distributions of the remaining features can be found in Appendix D.
Figure 5: Feature distributions of the original data and generated prediction-risky samples.
Parameter-sensitive samples. This experiment investigates data samples that are sensitive to small perturbations in the model parameters. Unlike prediction-risky samples, parameter-sensitive samples may exist anywhere in the input space, as their classification changes with slight shifts in the model’s parameters. To guide this analysis, we pose the following question:
What kind of data samples vary in classification due to small changes in model parameters?
We train an MLP on the FICO dataset and generate parameter-sensitive samples by perturbing the model parameters using a Gaussian distribution centered at the original weights with fixed variance. Using the probing function in (6), we generate and analyze 500 such samples to identify instances most susceptible to model variation, and compare them with prediction-risky samples.
Figure 6 shows density plots of four representative features (see Appendix D for additional ones), comparing parameter-sensitive and prediction-risky samples. By comparing these two distributions, we gain insights into how the model perceives uncertainty from different perspectives. While the prediction-risky samples are associated with uncertainty near the decision boundary, the parameter-sensitive samples highlight regions in the feature space where small parameter changes can flip predictions. The features AverageMinFile (average observation period) and NumTotalTrades (total number of trades) exhibit similar distributions across both sets. In contrast, the features MSinceMostRecentTradeOpen (months since most recent trade) and NumInqLast6M (inquiries in the past six months) diverge. For example, NumInqLast6M, which signals recent credit-seeking activity, is lower among prediction-risky samples, indicating that individuals with fewer recent inquiries are more likely to fall near the decision boundary. In contrast, parameter-sensitive samples exhibit a broader distribution, indicating that parameter shifts affect individuals across a wider range of credit inquiry patterns. This may be because frequent inquiries reflect diverse financial behaviors, making these samples more vulnerable to prediction instability. These findings suggest that some features contribute more to robustness under parameter variation, while others primarily influence boundary-sensitive classifications.
Figure 6: Feature distributions in generated parameter-sensitive and prediction-risky samples.
Fixed-label samples. We apply the probing function $G$ in (7) to the Adult dataset [2], a widely used benchmark for binary classification based on income level (whether an individual earns more than \$50K annually). A logistic regression model is trained on the dataset, and we examine its behavior by constructing counterfactual samples. Specifically, given a factual instance $({\boldsymbol{\mathbf{x}}}_{0},y)$, we generate samples using the probing function with $y^{\prime}\neq y$ and regularizer $R({\boldsymbol{\mathbf{x}}})=\|{\boldsymbol{\mathbf{x}}}-{\boldsymbol{\mathbf{x}}}_{0}\|^{2}$. In this experiment, the factual instance represents a Latin-American Black Female, predicted to earn less than \$50K. Using our framework, we pose the following question:
What feature changes would cause the model to predict an income above \$50K for this individual?
To address this question, the probing function is designed to balance two objectives: steering the model prediction toward the target label $y^{\prime}=1$ (using cross-entropy loss), and staying close to the original input (via the regularizer term $R_{G}({\boldsymbol{\mathbf{x}}})$).
Figure 7 shows the distribution of the generated samples. The results provide insights into the model’s classification process and the factors it deems influential in income predictions. While generating counterfactual samples, we impose bounds on age, educational attainment, and weekly working hours, which are enforced during Langevin dynamics sampling process by clipping each step to remain within the specified ranges. Additionally, all generated samples are projected to lie within the original feature ranges from the given training and test sets. In the figure, the factual instance is marked in red, with a vertical dashed line for numerical features and a red text label for categorical ones. Comparing the factual input with the counterfactual distribution reveals significant categorical shifts. For example, the majority of samples indicate a change in gender from female to male, and a region shift from Latin America to Western Europe, suggesting that these features significantly influence the model’s decision. These observations raise questions about fairness and bias. While we may directly investigate the logistic regression coefficients associated with these features, e.g., female $\approx-1.375$, male $\approx-1.243$, the bias is more clearly revealed through the generated samples. More importantly, such coefficients are not readily available for more complex models like deep networks.
Figure 7: Feature distributions of generated counterfactual samples (blue shaded) with factual instance highlighted (red markers).
5 Conclusion and Limitations
We introduce a mathematical framework to probe trained models with tailored data samples for specific queries, going beyond traditional interpretability methods. By formulating probing functions, we demonstrate how to generate samples for scenarios like prediction risky, parameter sensitivity, and model contrast. Our computational study shows the framework’s effectiveness in classification and regression tasks on diverse datasets, revealing insights into model decision boundaries and input sensitivity. Our goal is to understand machine learning models for positive societal impact and our tools aid in analyzing and comprehending trained models through sample generation.
Our framework has also limitations. First, scaling to larger models, particularly deep learning architectures with billions of parameters, poses computational challenges. The iterative optimization and sampling procedures may become prohibitively expensive in such settings. Furthermore, due to implicit constraints among the features, our method may generate samples that are not representative enough of the dataset, potentially leading to narrow conclusions.
A promising avenue for future research is to explore the incorporation of constraints among features in order to ensure that the generated samples are realistic and align with known data dependencies. One approach could involve enforcing domain-specific relationships, such as monotonic constraints among features, which could improve the interpretability and accuracy of the generated samples. An additional direction for further research could be exploring how a model’s internal structure, such as hidden layers or activation functions, can be used to guide the sampling process.
Appendix A Linear Regression with Gaussian Data
We start with $y_{\boldsymbol{\theta}}(\mathbf{x})=\boldsymbol{\theta}^{\top}\mathbf{x}$ and $\ell_{F}(y,y^{\prime})=\ell_{G}(y,y^{\prime})=\frac{1}{2}(y-y^{\prime})^{2}$. Given a dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, we construct the loss function $F({\boldsymbol{\theta}})$ as the integral of $\ell_{F}(y_{\boldsymbol{\theta}}(\mathbf{x}),y)$, over the data distribution, which is approximated by the Dirac delta comb $\nu=\frac{1}{N}\sum_{i=1}^{N}\delta_{(\mathbf{x}_{i},y_{i})}$:
$F({\boldsymbol{\theta}})=\int_{\mathcal{X}\times\mathcal{Y}}\!\!\!\!\ell_{F}(y_{\boldsymbol{\theta}}(\mathbf{x}),y)\mathrm{d}\,\nu(\mathbf{x},y)=\frac{1}{2N}\sum_{i=1}^{N}|\mathbf{x}_{i}^{\top}{\boldsymbol{\theta}}-y_{i}|^{2}.$
Assume, for convenience, that a constant feature of 1 is included as the last coordinate of $\mathbf{x}$, allowing us to explicitly represent the intercept. Using this notation, we define
$\mathbf{x}=\begin{bmatrix}\mathbf{f}\\ 1\end{bmatrix},\qquad{\boldsymbol{\theta}}=\begin{bmatrix}{\boldsymbol{\xi}}&b\end{bmatrix},\text{ so that }\mathbf{x}^{\top}{\boldsymbol{\theta}}=\mathbf{f}^{\top}{\boldsymbol{\xi}}+b.$
We write the design matrix as
$D=\left[\begin{array}[]{cccc}\cdots&\mathbf{x}_{1}^{\top}&\cdots&1\\ \cdots&\mathbf{x}_{2}^{\top}&\cdots&1\\ &\vdots&&\vdots\\ \cdots&\mathbf{x}_{N}^{\top}&\cdots&1\end{array}\right]=\left[\begin{array}[]{cc}X&\mathbf{1}\end{array}\right].$
The quadratic loss function can then be expressed as
$F({\boldsymbol{\theta}})=\frac{1}{2N}\|D{\boldsymbol{\theta}}-\mathbf{y}\|^{2},$
where $\mathbf{y}=\begin{bmatrix}y_{1}&y_{2}&\cdots&y_{N}\end{bmatrix}^{\top}$ is the label vector. We can reorder the terms so that
$F({\boldsymbol{\theta}})$ $=\frac{1}{2N}(D{\boldsymbol{\theta}}-y)^{\top}(D{\boldsymbol{\theta}}-y)=\frac{1}{2N}\left({\boldsymbol{\theta}}^{\top}D^{\top}D{\boldsymbol{\theta}}-2{\boldsymbol{\theta}}^{\top}X^{\top}\mathbf{y}\right)+\text{const.}$.
$=\frac{1}{2}({\boldsymbol{\theta}}-\widehat{{\boldsymbol{\theta}}})^{\top}\frac{D^{\top}D}{N}({\boldsymbol{\theta}}-\widehat{{\boldsymbol{\theta}}})+\text{const}.$
where $\widehat{{\boldsymbol{\theta}}}=(D^{\top}D)^{-1}D^{\top}\mathbf{y}$. Note that this is precisely the ordinary least squares solution.
Since the loss function is quadratic, we can explicitly write the Gibbs distribution (which is the unrestricted solution to the Bayesian Learning Problem with $F$) as the Gaussian distribution
$q^{*}(\theta)\propto e^{-\beta F(\theta)}\propto e^{-\frac{1}{2}({\boldsymbol{\theta}}-\widehat{{\boldsymbol{\theta}}})^{\top}\frac{D^{\top}D}{N/\beta}({\boldsymbol{\theta}}-\widehat{{\boldsymbol{\theta}}})}\quad\text{thus }q^{*}(\mathbf{\theta})=\mathcal{N}\left(\widehat{{\boldsymbol{\theta}}},\left(\frac{D^{\top}D}{N/\beta}\right)^{-1}\right).$
Here, the variable $\beta$ is the inverse temperature defined as $\beta=1/\tau$.
Next, we construct $G$, a loss function on $\mathcal{X}\times\mathcal{Y}$. By fixing the label, we may also consider $G$ as a loss function only on $\mathcal{X}$, from which we derive a distribution over $\mathcal{X}$. To avoid overusing $\mathbf{x}$ and $y$, we denote elements of the labeled dataset as $(\mathbf{z},w)\in\mathcal{X}\times\mathcal{Y}$ with $\mathbf{z}=\left[\begin{smallmatrix}\mathbf{f}\\ 1\end{smallmatrix}\right]$. Using the first and second moments of Gaussians, we calculate
$G(\mathbf{z},w)$ $=\int_{\Theta}|\mathbf{z}^{\top}{\boldsymbol{\theta}}-w|^{2}q^{*}({\boldsymbol{\theta}})\mathrm{d}\,\theta$
$=\mathbf{z}^{\top}\mathbb{E}_{q^{*}}[{\boldsymbol{\theta}}{\boldsymbol{\theta}}^{\top}]\mathbf{z}-2w\mathbf{z}^{\top}\mathbb{E}_{q^{*}}[{\boldsymbol{\theta}}]+\text{ const}$
$=\mathbf{z}^{\top}\left(\widehat{{\boldsymbol{\theta}}}\widehat{{\boldsymbol{\theta}}}^{\top}+\left(\frac{D^{\top}D}{N\tau}\right)^{-1}\right)\mathbf{z}-2w\mathbf{z}^{\top}\widehat{{\boldsymbol{\theta}}}+\text{ const}.$
which is again a quadratic function in $\mathbf{z}.$ Let us now write this quadratic in terms of $\mathbf{f}$. We write $\widehat{{\boldsymbol{\theta}}}=\left[\begin{smallmatrix}\widehat{{\boldsymbol{\xi}}}\\ \widehat{b}\end{smallmatrix}\right]$.
First, a quick calculation gives the block diagonal form
[…]
where $A=\frac{X^{\top}X}{N}-\overline{\mathbf{x}}\overline{\mathbf{x}}^{\top}$ is the Schur complement and $\overline{\mathbf{x}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{x}_{i}$ is the mean data vector.
We can write $G$ as a quadratic function of $\mathbf{f}$ (fixing $w$) as
$G_{w}(\mathbf{f})$ $=\mathbf{f}^{\top}\left(\tau A^{-1}+\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\right)\mathbf{f}-2\mathbf{f}^{\top}\left(\tau A^{-1}\overline{\mathbf{x}}-\widehat{{\boldsymbol{\xi}}}\hat{b}+w\hat{b}\right)+\text{const.}$.
$=(\mathbf{f}-\widehat{\mathbf{f}})\left(\tau A^{-1}+\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\right)(\mathbf{f}-\widehat{\mathbf{f}})+\text{const}.$
Here, $\widehat{\mathbf{f}}$ is calculated as
$\widehat{\mathbf{f}}$ $=\left(\tau A^{-1}+\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\right)^{-1}\left(\tau A^{-1}\overline{\mathbf{x}}+\widehat{{\boldsymbol{\xi}}}(w-\hat{b})\right)$
$=\left(A_{\tau}-\frac{A_{\tau}\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}}{1+\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}}\right)\left(A_{\tau}^{-1}\overline{\mathbf{x}}+\widehat{{\boldsymbol{\xi}}}^{\top}(w-\hat{b})\right),$
where $A_{\tau}=\frac{1}{\tau}A$ and the Sherman-Morrison formula is used for inverting the matrix.
Now expanding the product, we obtain
$\widehat{\mathbf{f}}=\overline{\mathbf{x}}+A_{\tau}\widehat{{\boldsymbol{\xi}}}(w-\hat{b})-\frac{A_{\tau}\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\overline{\mathbf{x}}}{1+\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}}-A_{\tau}\widehat{{\boldsymbol{\xi}}}\frac{\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}}{1+\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}}(w-\hat{b}).$
Note that if we denote the predictions of the linear model as $\mathbf{x}_{i}^{\top}\widehat{{\boldsymbol{\xi}}}+\widehat{b}=\hat{y}_{i}$, we can rewrite the above formula as follows:
$\widehat{\mathbf{f}}$ $=\bar{\mathbf{x}}+A_{\tau}\widehat{{\boldsymbol{\xi}}}(w-\widehat{b})\frac{1}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}-A_{\tau}\hat{{\boldsymbol{\xi}}}\frac{\hat{{\boldsymbol{\xi}}}^{\top}\bar{\mathbf{x}}}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}$
$=\bar{\mathbf{x}}+A_{\tau}\hat{{\boldsymbol{\xi}}}\frac{(w-\hat{b})}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}-A_{\tau}\hat{{\boldsymbol{\xi}}}\frac{(\widehat{\overline{\mathbf{y}}}-\hat{b})}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}$
$=\bar{\mathbf{x}}+A_{\tau}\hat{{\boldsymbol{\xi}}}\frac{w-\widehat{\overline{\mathbf{y}}}}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}.$
Here, we denoted the prediction of the average data by $\widehat{\overline{\mathbf{y}}}=\hat{\xi}^{\top}\bar{\mathbf{x}}=\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}$.
Finally, let’s rewrite $A_{\tau}\widehat{{\boldsymbol{\xi}}}$ and $\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}$ in terms of interpretable statistical quantities. Recall that $A_{\tau}=\frac{1}{\tau}\left(\frac{X^{\top}X}{N}-\overline{\mathbf{x}}\overline{\mathbf{x}}^{\top}\right)$. Using this, we compute
$A_{\tau}\widehat{{\boldsymbol{\xi}}}$ $=\frac{1}{\tau}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{x}_{i}(\underbrace{\mathbf{x}_{i}^{\top}\widehat{{\boldsymbol{\xi}}}}_{=\hat{y}_{i}-\hat{b}})-\overline{\mathbf{x}}\underbrace{\bar{\mathbf{x}}^{\top}\hat{{\boldsymbol{\xi}}}}_{=\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}-b}\right)$
$=\frac{1}{N\tau}\sum_{i=1}^{N}(\mathbf{x}_{i}-\overline{\mathbf{x}})(\hat{y}_{i}-\hat{b})$
$=\frac{1}{N\tau}\sum_{i=1}^{N}(\mathbf{x}_{i}-\overline{\mathbf{x}})(\hat{y}_{i}-\hat{\overline{\mathbf{y}}})$
$=\frac{1}{\tau}\operatorname{Cov}(X,\widehat{\mathbf{y}}).$
In the final expression, the term $\operatorname{Cov}(X,\widehat{\mathbf{y}})$ corresponds directly to the previous line. This covariance is a vector that averages data deviations, weighted by prediction deviations. In the line before last, we replaced $\widehat{b}$ with any constant since it is independent of $i$, and the first factor sums to the zero vector. Additionally, we leveraged a key property of linear models: the average of the predictions is the same as the prediction of the average.
A similar calculation yields,
$\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}=\frac{1}{\tau}\left(\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}^{2}-\left(\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}\right)^{2}\right)=\frac{1}{\tau}\operatorname{Var}(\widehat{\mathbf{y}}).$
Therefore, we obtain an explicit quadratic formulation of the data loss function $G$ in terms of $\mathbf{f}$ at a fixed $w$. This means that the data distribution $p^{*}(\mathbf{x})$, which solves the unrestricted Bayesian Learning Problem, follows a Gaussian distribution given as
$p^{*}(\mathbf{f})\propto e^{-G_{w}(\mathbf{f})}\propto\mathcal{N}(\widehat{\mathbf{f}},\Sigma),$
where
$\widehat{\mathbf{f}}=\overline{\mathbf{x}}+\frac{\operatorname{Cov}(X,\widehat{\mathbf{y}})}{\tau+\operatorname{Var}(\widehat{\mathbf{y}})}\left(w-\frac{1}{N}\sum_{i=1}^{N}\widehat{y}_{i}\right),$
and
$\Sigma^{-1}=\left(\tau\left(\frac{X^{\top}X}{N}-\bar{\mathbf{x}}\bar{\mathbf{x}}^{\top}\right)^{-1}+\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\right).$
The interpretation of the mean $\widehat{\mathbf{f}}$ is as follows: if you want to sample from a data distribution that will produce a given $\omega$, then you should not sample around $\bar{\mathbf{x}}$ (which would be the case without output restrictions). Instead, you shift $\bar{\mathbf{x}}$ in proportion to the difference between $\omega$ and the mean of the training label predictions, following the direction of the covariance between the training data and predicted labels.
Appendix B Computational Setup
In this section, we supplement our computational study by presenting the datasets and detailing the neural network architectures.
B.1 Datasets Used in the Experiments
Our experiments are conducted using three numerical datasets and one visual dataset from the literature. The details of the datasets are provided below.
Adult. The Adult dataset (CC BY 4.0 license), derived from the 1994 Census database, comprises 48,842 observations with 14 features, including both continuous and categorical variables [2]. The primary objective is to classify individuals based on whether their annual income exceeds \$50,000 USD. Data preprocessing steps are applied to address missing values and handle categorical features. We applied one-hot encoding to transform the categorical features into a numerical format suitable for our framework.
FICO. The FICO (HELOC) dataset (CC0: Public Domain) consists of home equity line of credit applications submitted by homeowners [7]. It includes 10,459 records with 23 features, comprising both numerical and ordinal variables. The primary objective is to classify applications based on their risk performance, identifying whether an applicant is likely to meet payment obligations or become delinquent. Data preprocessing steps are applied to address missing values.
Housing. The Housing dataset (CC0: Public Domain), sourced from Kaggle, includes information on various house attributes such as lot size, number of rooms, and number of stories [13]. The dataset contains 535 records and 12 features, comprising both numerical and ordinal variables. The primary objective is to predict housing prices based on these features.
MNIST. The MNIST dataset (Data files © Original Authors) is a widely used benchmark in computer vision, consisting of 70,000 grayscale images of handwritten digits (0–9), each represented as a 28×28 pixel matrix [16]. The dataset is divided into 60,000 training samples and 10,000 test samples. The primary objective is to classify images based on the digit they represent. We normalized each of the images to be arrays of shape $(28,28,1)$ with FP32 values in the interval $[0,1]$.
B.2 Experimental Setup
We ran all the experiments on Apple M1 Pro with 32 GB RAM.
For the parameter-sensitive and prediction-risky experiments on the FICO dataset, we trained an MLP with ReLU activation functions and layer widths of $128-32-8-2$. Dropout with a rate of 0.2 was applied after each activation layer to prevent overfitting. The model was trained using a batch size of $128$ for $10,000$ steps.
[…]
Both the CNN and MLP models were trained for 10,000 update steps using a batch size of 128 and the Adam optimizer. The learning rate followed an exponential decay schedule, starting with a maximum learning rate of 0.1, decaying by a rate of 0.9 every 100 steps.
Appendix C Langevin Dynamics
We sample data points in this paper by solving a Bayesian optimization problem over distributions on the data space:
$p^{*}=\operatorname*{arg\,min}_{p\in\mathcal{Q}}\mathbb{E}[G]-\tau\mathcal{H}(p).$
where $\mathcal{H}(p)$ denotes the entropy and $\tau>0$ is a temperature parameter. On unconstrained data space, i.e., when $\mathcal{Q}=\mathcal{P}(\mathbb{R}^{d})$, one approach of sampling from $p^{*}$ is to simulate the Langevin dynamics, which is a Stochastic Differential Equation (SDE) given by
$\frac{\mathrm{d}\,X}{\mathrm{d}\,t}=-\frac{1}{\tau}\nabla G(X)+\sqrt{2}\frac{\mathrm{d}\,W}{\mathrm{d}\,t}$
where $W$ is a Wiener process (i.e., standard Brownian motion), the $G$ is called a drift term. The limiting distribution of this SDE converges to $p_{\infty}(X)\propto e^{-\frac{1}{\tau}G(x)}$.
For concrete computation, we work with the standard discretization
$\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t}-\eta\nabla G(\mathbf{x}_{t})+\sqrt{2\eta\tau}\varepsilon_{t}$
where $\eta>0$ is the step size, and $\varepsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ is sampled independently from the unit normal distribution at each time step.
To improve convergence and ensure the correctness of the sampling distribution, we apply a Metropolis–Hastings acceptance step. Specifically, the proposed update $\mathbf{x}_{t+1}$ is accepted as $\mathbf{x}_{t+1}=\tilde{\mathbf{x}}_{t+1}$ with probability
$\alpha:=\min\left\{1,\frac{\pi(\tilde{\mathbf{x}}_{t+1})q(\mathbf{x}_{t}|\tilde{\mathbf{x}}_{t+1})}{\pi(\mathbf{x}_{t})q(\tilde{\mathbf{x}}_{t+1}|\mathbf{x}_{t})}\right\}$
or otherwise rejected, in which case $\mathbf{x}_{t+1}=\mathbf{x}_{t}$. Here, $\pi\propto e^{-\frac{1}{\tau}G}$ and $q(\mathbf{x}^{\prime}|\mathbf{x})$ is the transition probability of stepping from $\mathbf{x}$ to $\mathbf{x}^{\prime}$ given by $q(\mathbf{x}^{\prime}|\mathbf{x})=\exp(-\frac{1}{4\tau}\|\mathbf{x}^{\prime}-\mathbf{x}+\tau\nabla G(\mathbf{x})\|^{2})$. An implementation of this sampling procedure is provided in the submitted code repository, specifically in the file langevin.py using JAX.
Other Methods and Variants
Beyond the Langevin dynamics approach, several alternative methods exist for sampling data points. For instance, one may use the Picard iteration method or various splitting schemes (see [9]) to discretize and solve the above SDE. Another approach is to restrict the distribution family $\mathcal{Q}$ to be a tractable statistical manifold and apply Variational Inference techniques [8].
It is also worth noting that when sampling over a data manifold, if there exists a mapping $\varphi:\mathbb{R}^{d}\to X$ and $G:X\to\mathbb{R}$, then the corresponding Gibbs distribution on $\mathbb{R}^{d}$ (with respect to the Lebesgue measure) is given by $e^{-\frac{1}{\tau}H}$, where $H=G\circ\varphi:\mathbb{R}^{d}\to\mathbb{R}$. Samples drawn from this distribution can be pushed forward via $\varphi$ to obtain samples in $X$. These pushed-forward samples follow a distribution that can be interpreted as a Gibbs distribution over $X$ with respect to the base measure $\nu$, which is the pushforward of the Lebesgue measure under $\varphi$.
Appendix D Additional Numerical Results
This section presents additional results that complement the findings discussed in Section 4. These results provide further insights into the generated data distributions, feature variations, and model behavior under different probing scenarios.
Model-contrasting samples. This subsection extends our analysis of model-contrasting samples by applying the framework to a different tabular dataset. In this experiment, we examine prediction divergence between support vector regression (SVR) and linear regression (LR) models. We use the Housing dataset, where the primary objective is to predict house prices based on various structural and amenity-related features. We split the dataset into training-test sets and train both models on the same training data. To generate data samples where the two models diverge in their predictions, we formulate the cost function given in (5) as $\ell_{G}(y_{1},1-y_{2})=\exp(-(y_{1}-y_{2})^{2})$. Using our framework, we generate data samples to identify the regions of the input space where the models exhibit significant disagreement, likely due to their differing assumptions about feature interactions and predictive mechanisms.
Figure 8 presents a scatter plot comparing the predictions of the SVR and LR models. The blue points represent the predictions of the models in the test data, demonstrating that the two models generally produce highly similar outputs, with minimal differences observed. The green points, on the other hand, represent generated samples, highlighting instances where the models exhibit contrasting predictions. The zoomed-in inset further emphasizes these discrepant predictions, demonstrating that our framework effectively identifies and generates data points that maximize the divergence between the two models.
Figure 8: Comparison of SVR and LR predictions on test and generated data.
Figure 9: Features in test data and generated samples that produce different predictions for SVR and LR.
Figure 9 compares the feature distributions between the synthetic dataset generated by our framework and the test data. The box plots represent the range of values for each feature, with blue corresponding to the test data and green representing the generated samples. The Housing dataset [13] used in this experiment contains real estate information such as lot size, number of bathrooms, number of stories, and heating/air conditioning types, aiming to predict house prices. This figure provides a clear visualization of how the generated data differs from the test data in terms of feature distributions. For instance, as the number of bathrooms and stories increases, the model predictions diverge. Additionally, hot water heating and air conditioning exhibit a distinct concentration in the synthetic data, with most generated samples clustering around higher values compared to the test data. This suggests that these features play a prominent role in distinguishing instances where the models behave differently. Overall, this figure offers insights into how the generated samples differ from the original dataset, highlighting key feature distributions that drive divergence in model predictions and providing a deeper understanding of how our framework probes model behavior.
Prediction-risky samples. To further investigate data samples near the decision boundary, we present the distributions of all features in the original dataset and the generated prediction-risky samples in Figure 10. These density plots provide a comprehensive view of the differences between the generated samples and the original data across multiple features. By analyzing these distributions, we can observe how the model identifies borderline cases based on different financial attributes. Across multiple features, the generated prediction-risky samples exhibit a much narrower distribution compared to the original data. This suggests that the model focuses on a specific subset of feature values when identifying borderline cases.
Figure 10: Feature distributions in the original data and generated prediction-risky samples.
Parameter-sensitive samples.To complement the findings presented in Section 4, we provide the full set of feature distributions comparing parameter-sensitive samples and prediction-risky samples in Figure 11. These density plots illustrate how the two types of generated samples differ. By analyzing these distributions, we observe that while some features exhibit similar trends across both sample types, others show notable divergences. Features with broader distributions in parameter-sensitive samples indicate that model perturbations impact a wider range of instances.
Figure 11: Feature distributions in generated parameter-sensitive and prediction-risky samples.
Fixed-label samples. We now analyze a different factual instance from the original data to further investigate the model’s behavior. The factual instance considered represents a Latin American white male who is predicted to earn more than \$50K. To explore the conditions under which the model would classify this individual as earning less than \$50K, we generate a set of counterfactual samples. Figure 12 presents the distribution of these generated counterfactual samples, highlighting the key feature variations that lead to a different classification outcome. In the generated counterfactual samples, while no categorical changes are observed, the numerical features age, educational attainment, and working hours exhibit lower values compared to the factual instance, implying that a reduction in these features leads to a shift in classification.
Figure 12: Feature distributions of generated counterfactual samples (blue shaded) with factual instance highlighted (red markers).
Appendix E The use of VAEs
A notable example of using pushforwards to obtain points on the data manifold comes from image datasets. We employ a VAE architecture with two convolutional layers each for the encoder and decoder submodules. Features in the convolutional layers are $32$ and $64$ with kernel sizes of (3,3) and a stride of (2,2). During training, the reconstruction loss is computed using bitwise entropy.
Figure 13 shows how this setup works for constructing loss functions $G$ on the latent space. One may use a combination of models, each precomposed with the decoder of the trained VAE. The resulting distribution on the latent space, after pushforwarding (i.e., passing the samples through the decoder), corresponds to a distribution on the data that is closer to the original data distribution.
Figure 13: By precomposing with the decoder submodule of a trained neural network, we can define $G$ functions on the lower-dimensional latent space, while still leveraging networks designed for higher-dimensional image inputs.